# On the Bergman metric of symmetric spaces

### v1

@@ 2.3. Final remarks
In the proof of Theorem 1.2, by [8, Lemma 1], the properties of the Bergman kernel, and the fact that a pluripolar set has zero Lebesgue measure, the restriction of functions gives a Hilbert spaces isomorphism
$A^{2}(\widetilde{\Omega})\to A^{2}(\widetilde{\Omega}\smallsetminus E):\psi\mapsto\psi|_{\widetilde{\Omega}\smallsetminus E}$
In particular $K_{\widetilde{\Omega}\smallsetminus E}=K_{\widetilde{\Omega}}|_{(\widetilde{\Omega}\smallsetminus E)\times(\widetilde{\Omega}\smallsetminus E)}$, which implies $g_{\widetilde{\Omega}\smallsetminus E}=g_{\widetilde{\Omega}}|_{\widetilde{\Omega}\smallsetminus E}$, i.e. $\lambda=1$.
@@
The converse of Theorem 1.1 clearly holds, since the Bergman metric of a bounded symmetric domain (and in general, the metric of a symmetric space) is complete. A “converse” result of Theorem 1.2 is also true. To be precise, if $\widetilde{\Omega}\subset\mathbb{C}^{n}$ is a bounded symmetric domain and $E\subset\Omega$ is pluripolar and closed in $\Omega$, then $g_{\widetilde{\Omega}\smallsetminus E}=g_{\widetilde{\Omega}}|_{\widetilde{\Omega}\smallsetminus E}$ as observed in Remark 2.3, resulting into $\nabla R^{\widetilde{\Omega}\smallsetminus E}=0$.
@@
With regards to Theorem 1.2, we remark that the uniformization in [7] actually holds for a more general class of Kähler manifolds, that is, Stein manifolds such that their Bergman space
  • is nontrivial and base-point free;
  • separates points;
  • separates holomorphic dimensions.
This conditions ensure that the Bergman kernel and metric of the manifold are well-defined and that it is possible to define a Bergman-Bochner map, which turns out to be injective. Since a Stein Hermitian symmetric space is of noncompact type and the work of Calabi holds for real analytic Kähler manifolds, our methods can be applied to this more general setting. Thus, Theorem 1.1 and Theorem 1.2 hold for Stein manifolds that satisfy the above mentioned conditions.
Remark 2.6.
With regards to [5], it is natural to ask whether a parallel result is obtained by removing the assumption of pseudoconvexity of the domain. In Theorem 1.2 we proved, using only the local symmetry condition, that the bounded locally symmetric domain is biholomorphic to an open subset of a bounded symmetric domain with the property that the Bergman metric of such open subset is, after eventually rescaling, the Bergman metric of the bounded symmetric domain. Since the analogue of this fact in the constant holomorphic sectional curvature setting is the starting point in [5], we pose the following:
@@
Let $\Omega\subset\mathbb{C}^{n}$, $n\geq 1$, be a bounded domain. If $\nabla R^{\Omega}=0$, then there exist $\widetilde{\Omega}\subset\mathbb{C}^{n}$ bounded symmetric domain and $E\subset\widetilde{\Omega}$ of zero Lebesgue measure over which $A^{2}(\widetilde{\Omega}\smallsetminus E)$ functions extend holomorphically to $\widetilde{\Omega}$, such that $\Omega\cong\widetilde{\Omega}\smallsetminus E$.

### v2

@@ 2.3. Final remarks
Notice that the converse of Theorem 1.1 clearly holds, since the Bergman metric of a bounded symmetric domain (and, more generally, the metric of a symmetric space) is complete. We claim that an analogous “converse” statement for Theorem 1.2 is also true.
Indeed, in the proof of Theorem 1.2, by [8, Lemma 1], the basic properties of the Bergman kernel, and the fact that a pluripolar set has zero Lebesgue measure, the restriction map
$A^{2}(\widetilde{\Omega})\longrightarrow A^{2}(\widetilde{\Omega}\setminus E),\qquad\psi\longmapsto\psi|_{\widetilde{\Omega}\setminus E},$
is an isomorphism of Hilbert spaces. In particular, $K_{\widetilde{\Omega}\setminus E}=K_{\widetilde{\Omega}}\big|_{(\widetilde{\Omega}\setminus E)\times(\widetilde{\Omega}\setminus E)},$ and hence the associated Bergman metrics satisfy $g_{\widetilde{\Omega}\setminus E}=g_{\widetilde{\Omega}}\big|_{\widetilde{\Omega}\setminus E}.$
Therefore, if $\widetilde{\Omega}\subset\mathbb{C}^{n}$ is a bounded symmetric domain and $E\subset\widetilde{\Omega}$ is pluripolar and closed in $\widetilde{\Omega}$, then $(\widetilde{\Omega}\setminus E,g_{\widetilde{\Omega}\setminus E})$ inherits the same local symmetry: namely, since $g_{\widetilde{\Omega}\setminus E}$ coincides with the restriction of $g_{\widetilde{\Omega}}$, and the curvature tensor of the Bergman metric on a bounded symmetric domain is parallel, we obtain $\nabla R^{\widetilde{\Omega}\setminus E}=0,$ which proves the claim.
@@
Notice that the authors of [7] extend the uniformization theorem mentioned above for bounded pseudoconvex domains to Stein manifolds whose Bergman metric has constant holomorphic sectional curvature, provided that their Bergman space satisfies the following conditions:
  • it is nontrivial and base-point free;
  • it separates points;
  • it separates holomorphic directions.
These assumptions guarantee that the Bergman kernel and the Bergman metric are well-defined, and they allow one to define the Bergman–Bochner map, which in this setting is injective. Using similar techniques, one can extend Theorem 1.1 and Theorem 1.2 to Stein manifolds satisfying the above conditions and the local symmetry of the Bergman metric hypothesis.
@@
With regards to [5], it is natural to ask whether a parallel result can be obtained by removing the assumption of pseudoconvexity of the domain. In Theorem 1.2 we proved, using only the local symmetry condition, that the bounded locally symmetric domain is biholomorphic to an open subset of a bounded symmetric domain with the property that the Bergman metric of such open subset is, after eventually rescaling, the Bergman metric of the bounded symmetric domain. Since the analogue of this fact in the constant holomorphic sectional curvature setting is the starting point in [5], we pose the following:
@@
Let $\Omega\subset\mathbb{C}^{n}$, $n\geq 1$, be a bounded domain. If $\nabla R^{\Omega}=0$, then there exist $\widetilde{\Omega}\subset\mathbb{C}^{n}$ bounded symmetric domain and $E\subset\widetilde{\Omega}$ of zero Lebesgue measure over which $A^{2}(\widetilde{\Omega}\setminus E)$ functions extend holomorphically to $\widetilde{\Omega}$, such that $\Omega\cong\widetilde{\Omega}\setminus E$.